OpenStudy (anonymous):

Which of the following sets of numbers is a Pythagorean triplet? 16, 30, 32 16, 30, 34 24, 45, 48 24, 45, 54

Parth (parthkohli):

These are in the form a,b,c...now the only work you have to do is to check if \(a^2 + b^2 = c^2\)

Parth (parthkohli):

For example check if \(16^2 + 30^2 = 32^2\)

OpenStudy (maheshmeghwal9):

ya parthcohli:)

OpenStudy (anonymous):

I just did this (: 16, 30 and 34 :) Check that one :]

OpenStudy (paxpolaris):

you can divide by common factor and check ..........to make it easier

OpenStudy (paxpolaris):

8,15,16 8,15,17 8,15,16 8,15,18